# Vikhr: Constructing a State-of-the-art Bilingual Open-Source Instruction-Following Large Language Model for Russian

### v5

@@ Abstract
There has been a surge in the development of various Large Language Models (LLMs). However, text generation for languages other than English often faces significant challenges, including poor generation quality and reduced computational performance due to the disproportionate representation of tokens in the model’s vocabulary. In this work, we address these issues by developing a pipeline for adaptation of English-oriented pre-trained models to other languages and constructing efficient bilingual LLMs. Using this pipeline, we construct Vikhr, a state-of-the-art bilingual open-source instruction-following LLM designed specifically for the Russian language. “Vikhr” refers to the name of the Mistral LLM series and means a “strong gust of wind.” Unlike previous Russian-language models that typically rely on LoRA adapters on top of English-oriented models, sacrificing performance for lower training costs, Vikhr features an adapted tokenizer vocabulary and undergoes the continued pre-training and instruction tuning of all weights. This not only enhances the model’s performance but also significantly improves its computational and contextual efficiency. The remarkable performance of Vikhr across various Russian-language benchmarks can also be attributed to our efforts in expanding instruction datasets and corpora for continued pre-training. Vikhr not only sets the new state of the art among open-source LLMs for Russian but even outperforms some proprietary closed-source models on certain benchmarks. The model weights, instruction sets, and code are publicly available [^1] https://huggingface.co/Vikhrmodels.
@@ 1 Introduction
The aim of maximizing the LLM performance for a specific language within a certain number of parameters has led researchers to develop bilingual LLMs Sengupta et al. (2023); Pieri et al. (2024); Faysse et al. (2024). These LLMs prioritize a regional language, e.g. JaisSengupta et al. (2023) focuses on Arabic, but they are trained also on English data. The inclusion of English data in pre-training alongside regional language data is motivated by the significantly larger volume of English data available. This helps LLMs substantially enhance skills such as logical and common sense reasoning, which are also applied when generating text in a regional language. Bilingual LLMs is a perspective direction as they can remain small and efficient, but at the same time comprehensively capture linguistic nuances and cultural contexts of the regional language.
This work seeks to develop a pipeline for adapting English LLMs to other languages facilitating the development of bilingual LLMs. Specifically, we aim to build an instruction-following bilingual LLM for Russian and English that could be used for multilingual natural language processing research.
@@
Building even a small LLM tailored to a specific language from scratch demands a lot of computational resources. Consequently, many researchers opt to fine-tune LoRA adapters Hu et al. (2021) for English-oriented LLMs using some language-specific data. While this approach can improve model generation quality, it does not address computational inefficiency because the tokenizer and model vocabulary remain unchanged. In contrast, our approach not only fine-tunes a base LLM on Russian language data but also reconstructs its underlying tokenizer and vocabulary, alongside suggesting an improved method for continued pre-training. Additionally, we have significantly expanded the available Russian datasets for instruction tuning. The developed LLM achieves state-of-the-art results for the Russian language among other open-source counterparts across a wide range of benchmarks.
@@
  • We have developed a pipeline for adapting English-oriented LLMs to other languages. The pipeline implements vocabulary adaptation, continued pre-training with regularization to prevent “catastrophic forgetting”, and instruction tuning.
  • Using the pipeline, we have constructed Vikhr – a state-of-the-art open-source instruction-following LLM oriented on the Russian language. In addition to its high generation quality, Vikhr features an efficient tokenizer that enables rapid text generation and good context utilization.
@@
  • We conducted an extensive evaluation of several open-source LLMs on evaluation benchmarks for Russian, demonstrating that Vikhr achieves new state-of-the-art results. Ablation studies confirm the effectiveness and validity of the individual components within our LLM adaptation pipeline.
@@ 2 Related Work
One of the first prominent series of generative LLMs for Russian is ruGPT AI Forever (2022); Zmitrovich et al. (2023). The authors developed several models, trained on the standard language modeling task, with sizes reaching up to 13 billion parameters. These models were created from the scratch and trained on a large Russian corpus, enabling them to capture the linguistic nuances of Russian more effectively than multilingual models. Additionally, since the training data was mostly in Russian, these models also have efficient tokenization. However, the lack of multilingual data (e.g. in English) limits their performance. Notably, the ruGPT models are not instruction-tuned.
Gusev (2023) suggests to leverage reasoning capabilities of existing English-oriented LLMs and adapt them to the Russian language by training LoRA adapters. They created an Alpaca-like set of Russian instruction-output pairs and performed instruction tuning on it. As a result, they have established the series of models called Saiga, which demonstrate the competitive performance and used to be a reasonable choice for an off-the-shelf open-source Russian LLMs for the past years. However, the tokenizer in these models is not adapted, so they experience issues with the context size and computational efficiency.
Tikhomirov and Chernyshev (2023) address these issues in Saiga. In addition to model tuning on Russian data, they also adapted the model tokenizer. They note that improving tokenization helps to both enhance the efficiency of the model and its performance while reducing memory consumption. However, during continued pre-training, the authors freezed the model weights, except for LM heads and token embeddings, which leads to the suboptimal performance.
In this work, we take advantage of pre-trained English-oriented LLMs, adapt LLM tokenizer for better computational and contextual efficiency, leverage continued pre-training on vast Russian-language corpora with regularization for preventing “catastrophic forgetting”, construct a novel extended set of Russian instruction-output pairs, and perform instruction tuning. The created LLM adaptation pipeline along with the data for continued pre-training and instruction tuning enables Vikhr to achieve new state-of-the-art results for Russian, maintain high performance for English, and demonstrate high computational efficiency.
@@ 3 LLM Construction Pipeline
The construction of Vikhr starts from one of the English-oriented LLMs. In this work, we discuss the Vikhr model based on Mistral 7B Jiang et al. (2023). The strong logical and common-sense reasoning capabilities, as well as the extensive world knowledge present in Mistral LLMs provide an excellent starting point for our model. These features partially transfer to Vikhr, enhancing its performance in generating text in Russian. The process of the LLM adaptation to Russian starts with the vocabulary adaptation. Then we perform continued pre-training of the LLM on large Russian datasets to mitigate the vocabulary shift and introduce culture specific knowledge. Finally, we perform fine-tuning of Vikhr on a set of instruction-output pairs in Russian.
@@ 3.2 Continued Pre-training
The new vocabulary requires new embedding matrices and LM heads. The tokens that were present in the original vocabulary are initialized with the old embeddings, the new tokens are initialized by averaging the embeddings of their pieces in the original embedding matrix Hewitt (2021). The similar approach is also applied to LM heads. Training models with these modifications demands much more computational resources than the mainstream method of adapting LLMs to new languages using LoRA adapters Hu et al. (2021). This is because it involves continued pre-training of the entire model and requires much more language-specific data to mitigate the shift in the vocabulary.
The dataset for continued pre-training is constructed from high-quality sources, including Russian Wikipedia, news articles, scientific papers from peer-reviewed journals and conferences, and top 100k up-voted posts on Habr – a popular online blog community focused on technology, software development, and science. The statistics of these datasets is presented in Table 2. We performed deduplication of the collection on the level of paragraphs using the MIHash algorithm Cakir et al. (2017). Furthermore, we performed filtration of the collected data. For this purpose, we annotated 20k documents using GPT-4-turbo with the aim to assess their informativeness, usefulness for studying, grammatical correctness, style, and safety. Using these annotations, we fine-tuned a RuBERT-tinyDale (2021) filtration model and applied it to the deduplicated corpus. After filtration, the total number of tokens left for continued pre-training is 11 billion.
@@
To speed up the process of continued pre-training, we use an optimized Flash attention implementation [^2] https://huggingface.co/docs/optimum/bettertransformer/tutorials/convert. As an optimization algorithm, we leverage AdamW, as it trades some memory efficiency in favor of robustness to the hyperparameter choice. The hyperparameters used for continued pre-training are presented in Table 3.
@@ 3.3 Instruction Tuning
Instruction tuning is an essential step in reaching high zero-shot performance with LLMs. It also allows to obtain a more natural communication with the model without complex prompting. Further fine-tuning techniques such as RLHF Ouyang et al. (2022) or DPO Rafailov et al. (2024), which require input from the assessors, are also crucial for such tasks as multicriteria alignment. However, the most significant performance gains are still achieved through instruction tuning Jha et al. (2023).
Previously, Gusev (2023) constructed an open-source set of instruction-output pairs for the Russian language (Saiga). The core Saiga dataset was created similar to Alpaca by querying ChatGPT (gpt-3.5-turbo) Taori et al. (2023). In this work, we extend this set by translating two English instruction datasets. First, we translated instructions for the FLAN model Wei et al. (2021) and generated answers in Russian using ChatGPT. Originally, FLAN instructions were constructed automatically from annotated datasets using templates to facilitate multitask and zero-shot capabilities of seq2seq models. Later, it was shown that this data also helps to improve decoder-only chat-oriented models as well. Second, we construct Veles [^3] https://huggingface.co/datasets/Vikhrmodels/Veles-2.5 by translating the English OpenHermes Teknium (2023) instruction dataset. Third, we incroporate without translation Nectar [^4] https://huggingface.co/datasets/berkeley-nest/Nectar Zhu et al. (2023) – an English instruction dataset. This ensures that Vikhr maintains strong performance in English as well.
Similar to the corpus for continued pre-training, we performed deduplication of the instruction set. Since the majority of the outputs were machine generated there are many low quality outputs. To mitigate this problem, we filtered out low quality pairs using a reward model trained on human data. For the reward model, we selected the multilingual-e5-large model Wang et al. (2024a). This model was particularly suitable for our needs due to its ability to handle multilingual data efficiently, ensuring that the classifier could accurately assess the quality of responses in both Russian and English. We trained the reward model on the answer preference dataset [^5] https://huggingface.co/datasets/Vikhrmodels/sbs, which was collected from human-written prompts and annotated using GPT-4. By applying this reward model, we filtered out low-quality instruction-output pairs, significantly enhancing the overall performance and reliability of our instruction datasets. The statistics of the Vikhr instruction datasets is presented in Table 4.
@@
LLM
Pre-train on
Russian
Training
Method
En-MMLU Ru-MMLU
Ru-MMLU
pro
CheGeKa
Russian
SuperGLUE
MERA
MTS AI Chat 7B (closed-source) $\diamondsuit$ false sft+dpo - 0.69 - 0.083 0.56 0.479
GigaChat-7B (closed-source) $\diamondsuit$ true sft+dpo - 0.67 - 0.451* 0.71* 0.479
aya101 false pt+sft 0.41 0.37 0.12 0.005 0.36 0.320
Mistral-7B-Instruct-v0.2 false none 0.60 0.78 0.11 0.005 0.57 0.400
rccmsu/ruadapt-mistral-7b-v0.1 false pt+sft 0.61 0.72 0.16 0.005 0.64 0.421
rugpt13b true none 0.25 0.25 0.10 0.132 0.52 0.208
saiga-mistral-7b-lora false sft 0.60 0.76 0.14 0.223 0.64 0.442
saiga-llama3-8b false sft 0.59 0.78 0.22 0.225 0.66 0.476
(Ours) Vikhr-7B-instruct_0.2 true pt+sft 0.62 0.80 0.19 0.231 0.67 0.485
Table 6: Evaluation results for Russian and multilingual LLMs. Pre-train on Russian means that the model underwent (continued) pre-training on Russian data. The following abbreviations are used: sft – instruction tuning, pt – (continued) pre-training; dpo – direct preference optimization. $\diamondsuit$ The results for GigaChat and MTS AI are taken from the leaderboards. The best result among open-source models is highlighted with bold, the second best is underscored. The best result among closed-source proprietary models is marked with *.
@@
Contrary to Saiga, we do not use LoRA adapters and just as in the phase of continued pre-training, we update all model parameters. The hyperparameters for the instruction tuning phase are presented in Table 5.
@@ 3.4 Hardware
Vikhr was trained on eight NVIDIA A100 GPUs 80GB. We spend approximately 1,000 GPU hours for the continued pre-training phase and 60 hours for instruction tuning.
@@ 4.1 Experimental Setup
Benchmarks.
The evaluation was performed on MMLU (En-MMLU) Hendrycks et al. (2021), Ru-MMLU [^6] https://huggingface.co/datasets/NLPCoreTeam/mmlu_ru, Ru-MMLU-pro [^7] https://huggingface.co/datasets/Vikhrmodels/mmlupro-ru, ruXNLI Conneau et al. (2018), CheGeKa Mikhalkova and Khlyupin (2022), Russian SuperGLUE shavrina2020russiansuperglue, and MERA fenogenova2024mera. The MMLU (En-MMLU) benchmark assesses LLMs across 57 subjects through multiple-choice questions, measuring a model’s general knowledge and reasoning abilities. We utilize this benchmark to verify that the model retains its bilingual proficiency. For this dataset, we report the accuracy@1 score. We constructed Ru-MMLU and Ru-MMLU-pro by automatically translating original English MMLU and MMLU-pro to Russian. Translation was done by GPT-3.5 and GPT-4 respectively. Just as for MMLU, for these datasets, we report the accuracy@1 score. CheGeKa is based on the questions from the Russian version of “Jeopardy” (the “Own game”). This benchmark contains challenging open-ended questions, requiring logical reasoning and world knowledge. It includes 29,376 training and 416 test instances. The reported evaluation metric is the F1 score. Russian SuperGLUE is a benchmark similar to well-known English SuperGLUE Wang et al. (2019). It tests LLMs on various natural language understanding tasks like reading comprehension and textual entailment. The metric reported in the results is accuracy@1. The MERA benchmark encompasses 21 evaluation tasks for generative LLMs in 11 skill domains. Note that among other tasks, MERA also includes CheGeKa, RuMMLU, and one of the subtasks of SuperGLUE (RWSD). The reported evaluation metric is the total score, which is the average of scores across all non-diagnostic tasks.
Baselines.
We compare Vikhr to six open-source and two proprietary closed-source competitors of the similar size. Open-source models: aya101 Üstün et al. (2024) – a massively multilingual LLM from CohereAI that follows instructions in 101 languages [^8] https://huggingface.co/CohereForAI/aya-101, it shows state-of-the-art results among massively multilingual LLMs; Mistral-7B-0.2-instruct Jiang et al. (2023) – an English-oriented LLM that was used as the base model for Vikhr; rccmsu/ruadapt_mistral_saiga_7b_v0.1 Tikhomirov and Chernyshev (2023) – a Russian-oriented LLM that was constructed from the Mistral model using similar adaptations of the tokenizer, token embeddings, and the LM head; saiga-mistral-7b-lora and saiga-llama3-8b Gusev (2023) – two versions of the Saiga models based on English-oriented LLMs and obtained by fine-tuning LoRA adapters on the Saiga instruction dataset [^9] https://huggingface.co/collections/IlyaGusev. Closed-source proprietary models for Russian: MTS AI Chat [^10] https://huggingface.co/MTSAIR/multi_verse_model and GigaChat-7b. The access to GigaChat weights is closed, so the reported results are taken from the leaderboards [^11] https://mera.a-ai.ru/ru/submits/10257. The results of MTS AI Chat are also taken from the leaderboard [^12] https://mera.a-ai.ru/ru/submits/10290.
4.2 Results in NLU Tasks
The evaluation results are presented in Table 6. As we can see, Vikhr outperforms all open-source models, including the ones that were built specifically for Russian. It also slightly outperforms its parent model Mistral on the En-MMLU benchmark, which might be the result of longer pre-training. The second place with close scores for all 4 Russian language benchmarks is obtained by the Saiga model based on recently released Llama-3. The high scores of this model probably are the result of the transfer of the outstanding performance of Llama-3. Since Saiga based on Llama-3 outperforms Saiga based on Mistral, we expect that applying our adaptation pipeline to Llama-3 would also help further improving the state of the art.
We note that the original Mistral-7B-0.2-instruct, despite being an English-oriented model, demonstrates competitive performance in 3 out of 4 Russian benchmarks. This demonstrates that such models could be viable alternatives at least for NLU tasks. The only dataset, where its performance is very low is CheGeKa, which is related to open-ended question-answering. This may be due to the lack of culture-specific knowledge, as the English-oriented model has not seen much Russian texts. Note that the MTS AI Chat also shows very low results on CheGeKa, which might also indicate the lack of culture-specific knowledge.
The proprietary model GigaChat substantially outperforms Vikhr on CheGeKa and notably on Russian SuperGLUE. We assume this is due to the use of much larger Russian datasets for pre-training. However, surprisingly, it falls behind Vikhr on Ru-MMLU. On all benchmarks, Vikhr outperforms the proprietary competitor from MTS AI.

### v6

# Vikhr: The Family of Open-Source Instruction-Tuned Large Language Models for Russian

@@ Abstract
There has been a surge in the development of various Large Language Models (LLMs). However, text generation for languages other than English often faces significant challenges, including poor generation quality and reduced computational performance due to the disproportionate representation of tokens in the model’s vocabulary. In this work, we address these issues by developing a pipeline for the adaptation of English-oriented pre-trained models to other languages and constructing efficient bilingual LLMs. Using this pipeline, we construct Vikhr, a series of bilingual open-source instruction-following LLMs designed specifically for the Russian language. “Vikhr” refers to the name of the Mistral LLM series and means a “strong gust of wind.” Unlike previous Russian-language models that typically rely on LoRA adapters on top of English-oriented models, sacrificing performance for lower training costs, Vikhr features an adapted tokenizer vocabulary and undergoes the continued pre-training and instruction tuning of all weights. This not only enhances the model’s performance but also significantly improves its computational and contextual efficiency. We also expanded the instruction datasets and corpora for continued pre-training. The model weights, instruction sets, and code are publicly available. [^1] https://huggingface.co/Vikhrmodels
@@ 1 Introduction
The aim of maximizing the LLM performance for a specific language within a certain number of parameters has led researchers to develop bilingual LLMs Sengupta et al. (2023); Pieri et al. (2024); Faysse et al. (2024). These LLMs prioritize a regional language, e.g. Jais Sengupta et al. (2023) focuses on Arabic, but they are trained also on English data. The inclusion of English data in pre-training alongside regional language data is motivated by the significantly larger volume of English data available. This helps LLMs substantially enhance skills such as logical and common sense reasoning, which are also applied when generating text in a regional language. Bilingual LLMs are a promising direction as they can remain small and efficient, but at the same time comprehensively capture linguistic nuances and cultural contexts of the regional language.
This work seeks to develop a pipeline for adapting English LLMs to other languages, facilitating the development of bilingual LLMs. Specifically, we aim to build an instruction-following bilingual LLM for Russian and English that could be used for multilingual natural language processing research.
@@
Building even a small LLM tailored to a specific language from scratch demands a lot of computational resources. Consequently, many researchers opt to fine-tune LoRA adapters Hu et al. (2021) for English-oriented LLMs using some language-specific data. While this approach can improve model generation quality, it does not address computational inefficiency because the tokenizer and model vocabulary remain unchanged. In contrast, our approach not only fine-tunes a base LLM on Russian language data but also reconstructs its underlying tokenizer and vocabulary, alongside suggesting an improved method for continued pre-training. Additionally, we have significantly expanded the available Russian datasets for instruction tuning.
@@
  • We have developed a pipeline for adapting English-oriented LLMs to other languages. The pipeline implements vocabulary adaptation, continued pre-training with regularization to prevent “catastrophic forgetting,” and instruction tuning.
  • Using the pipeline, we have constructed Vikhr – an open-source instruction-following LLM oriented on the Russian language. In addition to its high generation quality, Vikhr features an efficient tokenizer that enables rapid text generation and good context utilization.
@@
  • We conducted ablation studies of various pipeline components. The studies confirm the effectiveness and validity of the individual components within our LLM adaptation pipeline.
@@ 2 Related Work
One of the first prominent series of generative LLMs for Russian is ruGPT AI Forever (2022); Zmitrovich et al. (2023). The authors developed several models, trained on the standard language modeling task, with sizes reaching up to 13 billion parameters. These models were created from scratch and trained on a large Russian corpus, enabling them to capture the linguistic nuances of Russian more effectively than multilingual models. Additionally, since the training data was mostly in Russian, these models also have efficient tokenization. However, the lack of multilingual data (e.g. in English) limits their performance. Notably, the ruGPT models are not instruction-tuned.
Gusev (2023) suggests to leverage the reasoning capabilities of existing English-oriented LLMs and adapt them to the Russian language by training LoRA adapters. They created an Alpaca-like set of Russian instruction-output pairs and performed instruction tuning on it. As a result, they have established a series of models called Saiga, which demonstrate competitive performance and have been a reasonable choice for off-the-shelf open-source Russian LLMs for the past years. However, the tokenizer in these models is not adapted, so they experience issues with context size and computational efficiency.
Tikhomirov and Chernyshev (2023) address these issues in Saiga. In addition to model tuning on Russian data, they also adapted the model tokenizer. They note that improving tokenization helps to both enhance the efficiency of the model and its performance while reducing memory consumption. However, during continued pre-training, the authors froze the model weights, except for LM heads and token embeddings, which leads to suboptimal performance.
In this work, we take advantage of pre-trained English-oriented LLMs, adapt LLM tokenizer for better computational and contextual efficiency, leverage continued pre-training on vast Russian-language corpora with regularization for preventing “catastrophic forgetting”, construct a novel extended set of Russian instruction-output pairs, and perform instruction tuning. The created LLM adaptation pipeline along with the data for continued pre-training and instruction tuning enables Vikhr to achieve good generation quality for Russian and demonstrate high computational efficiency.
@@ 3 LLM Construction Pipeline
The construction of Vikhr starts from one of the English-oriented LLMs. In this work, we discuss the Vikhr model based on Mistral 7B Jiang et al. (2023). The strong logical and common-sense reasoning capabilities, as well as the extensive world knowledge present in Mistral LLMs, provide an excellent starting point for our model. These features partially transfer to Vikhr, enhancing its performance in generating text in Russian. The process of the LLM adaptation to Russian starts with the vocabulary adaptation. Then we perform continued pre-training of the LLM on large Russian datasets to mitigate the vocabulary shift and introduce culture-specific knowledge. Finally, we perform fine-tuning of Vikhr on a set of instruction-output pairs in Russian.
@@ 3.2 Continued Pre-training
The new vocabulary requires new embedding matrices and LM heads. The tokens that were present in the original vocabulary are initialized with the old embeddings; the new tokens are initialized by averaging the embeddings of their pieces in the original embedding matrix Hewitt (2021). The similar approach is also applied to LM heads. Training models with these modifications demands much more computational resources than the mainstream method of adapting LLMs to new languages using LoRA adapters Hu et al. (2021). This is because it involves continued pre-training of the entire model and requires much more language-specific data to mitigate the shift in the vocabulary.
The dataset for continued pre-training is constructed from high-quality sources, including Russian Wikipedia, news articles, scientific papers from peer-reviewed journals and conferences, and top 100k up-voted posts on Habr – a popular online blog community focused on technology, software development, and science. The statistics of these datasets are presented in Table 2. We performed deduplication of the collection on the level of paragraphs using the MIHash algorithm Cakir et al. (2017). Furthermore, we performed filtration of the collected data. For this purpose, we annotated 20k documents using GPT-4-turbo with the aim to assess their informativeness, usefulness for studying, grammatical correctness, style, and safety. Using these annotations, we fine-tuned a RuBERT-tinyDale (2021) filtration model and applied it to the deduplicated corpus. After filtration, the total number of tokens left for continued pre-training is 11 billion.
@@
To speed up the process of continued pre-training, we use an optimized Flash attention implementation. [^2] https://huggingface.co/docs/optimum/bettertransformer/tutorials/convert As an optimization algorithm, we leverage AdamW, as it trades some memory efficiency in favor of robustness to the hyperparameter choice. The hyperparameters used for continued pre-training are presented in Table 3.
@@ 3.3 Instruction Tuning
Instruction tuning is an essential step in reaching high zero-shot performance with LLMs. It also allows obtaining a more natural communication with the model without complex prompting. Further fine-tuning techniques such as RLHF Ouyang et al. (2022) or DPO Rafailov et al. (2024), which require input from the assessors, are also crucial for such tasks as multicriteria alignment. However, the most significant performance gains are still achieved through instruction tuning Jha et al. (2023).
Previously, Gusev (2023) constructed an open-source set of instruction-output pairs for the Russian language (Saiga). The core Saiga dataset was created similar to Alpaca by querying ChatGPT (gpt-3.5-turbo) Taori et al. (2023). In this work, we extend this set by translating two English instruction datasets. First, we translated instructions for the FLAN model Wei et al. (2021) and generated answers in Russian using ChatGPT. Originally, FLAN instructions were constructed automatically from annotated datasets using templates to facilitate multitask and zero-shot capabilities of seq2seq models. Later, it was shown that this data also helps to improve decoder-only chat-oriented models as well. Second, we construct Veles [^3] https://huggingface.co/datasets/Vikhrmodels/Veles-2.5 by translating the English OpenHermes Teknium (2023) instruction dataset. Third, we incorporate without translation Nectar [^4] https://huggingface.co/datasets/berkeley-nest/Nectar Zhu et al. (2023) – an English instruction dataset. This ensures that Vikhr maintains strong performance in English as well.
Similar to the corpus for continued pre-training, we performed deduplication of the instruction set. Since the majority of the outputs were machine generated, there are many low-quality outputs. To mitigate this problem, we filtered out low-quality pairs using a reward model trained on human data. For the reward model, we selected the multilingual-e5-large model Wang et al. (2024a). This model was particularly suitable for our needs due to its ability to handle multilingual data efficiently, ensuring that the classifier could accurately assess the quality of responses in both Russian and English. We trained the reward model on the answer preference dataset [^5] https://huggingface.co/datasets/Vikhrmodels/sbs, which was collected from human-written prompts and annotated using GPT-4. By applying this reward model, we filtered out low-quality instruction-output pairs, significantly enhancing the overall performance and reliability of our instruction datasets. The statistics of the Vikhr instruction datasets are presented in Table 4.
@@
Contrary to Saiga, we do not use LoRA adapters, and just as in the phase of continued pre-training, we update all model parameters. The hyperparameters for the instruction tuning phase are presented in Table 5.
@@ 3.4 Hardware
Vikhr was trained on eight NVIDIA A100 GPUs 80GB. We spent approximately 1,000 GPU hours for the continued pre-training phase and 60 hours for instruction tuning.
@@ 4.1 Experimental Setup
For evaluation, we use ruXNLI Conneau et al. (2018), and we also constructed Ru-MMLU and Ru-MMLU-pro by automatically translating the original English MMLU and MMLU-pro to Russian. Translation was done by GPT-3.5 and GPT-4, respectively. We report the accuracy@1 score.